OpenStudy (anonymous):

Find four consecutive even integers such that twice the sum of the second and third exceeds 3 times the first by 34.

OpenStudy (anonymous):

let first be x then (x+2+x+2)*2 = 3x + 34 2(2x+6) = 3x + 34 4x+12 = 3x + 34 x = 22 the numbers are 22 24 26 28